DEVELOPMENT OF A HIGHER RESOLUTION TERAHERTZ TIME-DOMAIN SPECTROMETER

\underline{DANIEL B. HOLLAND}, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125 (email to D.B.H.: holland@caltech.edu); GEOFFREY A. BLAKE, Division of Geological and Planetary Sciences, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125.

We report on our development of a terahertz (THz) time-domain spectrometer employing asynchronous optical sampling (ASOPS) between two commercial Ti:Sapphire lasers operating with a repetition rate of approximately 80 MHz. THz time-domain spectrometers typically achieve control of sampling time delay via the use of mechanical delay lines, yielding a practically feasible resolution limit of several GHz. There are commercially available THz time-domain spectrometers employing ASOPS with GHz repetition rate lasers; they achieve 1 GHz resolution. The use of our lower repetition rate lasers, while introducing technical challenges, offers a theoretical resolution limit of 80 MHz, an important benefit for our primary application of collecting THz spectra in support of astronomy in the far-infrared. Technical data on instrument design and performance as well as initial spectra are to be presented.